OpenStudy (lxelle):

A random variable X has probability density function given by f(x) = 2/3 x, 1 ≤ x ≤ 2, 0, otherwise. (i) Find E(X) (ii) Find P(X < E(X)) (iii) Hence explain whether the mean of X is less than, equal to or greater than the median of X. Workings and explanation please

OpenStudy (anonymous):

if i recall correctly \[E(x)=\int_{-\infty}^{\infty}xf(x)dx\] which in your case would be \[\int_1^2 \frac{2}{3}x^2dx\] get a second opinion though

OpenStudy (lxelle):

Yes bingo. I actually have difficulties in the last question.

OpenStudy (anonymous):

Refer to satellite's integral for \(E(X)\). I'll call the result \(k\). Second part wants the probability that \(X\) is less than the expectation \[P(X<k)=F(k)=\int_{-\infty}^kf(x)~dx=\int_1^k\frac{2}{3}x~dx\] where \(f\) denote the probability density function and \(F\) denotes the cumulative distribution function. The median is the value \(c\) that makes \[\int_{-\infty}^cf(x)~dx=\int_1^c\frac{2}{3}x~dx=\frac{1}{2}=n\] Compare \(n\) and \(k\).

OpenStudy (lxelle):

as for iii), in the answer scheme it says 115/243 < 1/2

OpenStudy (anonymous):

What are you getting for \(E(X)=k\)?

OpenStudy (lxelle):

14/9

OpenStudy (anonymous):

Okay, good. What's the value of \[\int_1^{14/9}\frac{2}{3}x~dx~~?\]

OpenStudy (lxelle):

115/243

OpenStudy (lxelle):

I didnt have any problem in i) and ii). I just need some clarification in iii). why is it 115/243<1/2 and not 14/9<1/2

OpenStudy (anonymous):

Right again. So you now have to find \(c\)... I just noticed that I included \(n\) for some reason... Ignore that. For (iii), you want to compare the values of mean \(k\) and the median \(c\). I don't see how saying \(\dfrac{115}{243}<\dfrac{1}{2}\) tells you anything about what you want to show... Here's how you compute the median: \[\int_1^c\frac{2}{3}x~dx=\frac{1}{2}~~\iff~~\left[\frac{x^2}{3}\right]_1^c=\frac{1}{2}~~\iff~~\frac{c^2}{3}-\frac{1^2}{3}=\frac{1}{2}\] which gives \(c=\sqrt{\dfrac{5}{2}}\approx1.581\) as the median. \(k=\dfrac{14}{9}\approx1.555\), so clearly the median is greater than the mean.

OpenStudy (lxelle):

I dont see it also. Anyways, thanks. Ill get back with my tutor. :)